OpenStudy (anonymous):

Having a brainfart xD Two rockets having the same acceleration start from rest, but the first rocket travels for twice as much time as the second one. If rocket A goes for 310 km, how far will B go? If rocket A reaches a speed of 320 , what speed will rocket B reach?

OpenStudy (anonymous):

v=v_0+at x-x_0=v_0t+0.5at^2 We know v_0=0, x_0=0, a=a (constant acceleration, same for both rockets), x=310 (for the first rocket), and v=320 (for the first rocket). So, let's say we half the time traveled. (1/2)^2=1/4, which is 1/4 the distance. (1/2)=(1/2), which is 1/2 the velocity. Do you understand?

OpenStudy (anonymous):

I think so. I was just confused at first because it only gave me the distance but not the time :p